A 10,000kg railroad car is travelling at 1.5m/s when it strikes another 9,000kg railroad car that is at rest. If the cars lock together, what is the final speed of the two railroad cars?


1
Expert's answer
2021-04-26T16:59:12-0400

The final speed can be found by momentum conservation principle:


"Mv=(m+M)u,\\\space\\
u=\frac{Mv}{m+M}=0.79\text{ m/s}."